OpenStudy (celinegirl):

check my answer - A The expression below shows the population of squirrels in a forest after 3x years: f(x) = 30(1.03)3x Which of the following is an equivalent function? f(x) = 30(3.0900)x f(x) = 30(1.0927)x f(x) = 10(1.0927)x f(x) = 10(3.0900)x

OpenStudy (boldjon):

f(x)=30(1.03)3^x or f(x)=30(1.03)3^x or f(x)=30(1.03)3^x?

OpenStudy (boldjon):

(a^n)m⟺a^n⋅^m f(x)=30(1.03)^3x⟹f(x)=30(1.03^3)^x⟹?

OpenStudy (boldjon):

f(x)=30(1.03)^3x⟹f(x)=30(1.03^3)^x⟹f(x)=30(1.092727)^x

OpenStudy (celinegirl):

Ohh

OpenStudy (boldjon):

and you can round that up to only say 1.0927

OpenStudy (mathmale):

Your f(x) = 30(1.03)3x really should be written as\[f(x) = 30(1.03)^{3x}\] via Equation Editor. If you're not comfortable with Equation Editor, please draw your equation so that you can show that 3x as the exponent it is.

OpenStudy (mathmale):

Following boldjon's comment: You can write \[f(x) = 30(1.03)^{3x}\]

OpenStudy (mathmale):

as\[f(x) = 30 *[(1.03)^{3} ]^x\]

OpenStudy (boldjon):

i don't know how to use the equation thing well @mathmale

OpenStudy (celinegirl):

so the answer is B?

OpenStudy (mathmale):

I just don't respond to "is it B?" questions. I'd be very happy to respond if y ou show your work and explain how and why you chose Option B.